# Balancing Efficiency and Expressiveness: Subgraph GNNs with Walk-Based Centrality

### v1

@@ E.1 Comparison to other centrality measures
To further assess the benefits of our specific centrality encoding for sampling subgraphs, we compared against using other centrality measures to sample subgraphs in the counting substructure task. We used Closeness centrality, Betweeness centrality, Pagerank centrality and Degree centrality as baselines using the Networkx library (Hagberg et al., 2008). From Fig. 6, we can see that using any of the different centrality methods performs better than random sampling across all substructures and number of samples. We also find that the Subgraph Centrality which we use, outperforms all other approaches in counting $3$ and $4$-cycles for any number of samples and in counting $3$ and $4$-paths when number of samples $\geq 5$. For $4$-cycles and other substructures, we find that Subgraph centrality is best, followed by Degree and Pagerank centrality and then Closeness and Betweenness centralities perform the worst of these centrality measures. This ranking of performance is aligned with how correlated these substructures are with the Subgaph Centrality on these synthetic graphs (as shown in Table 6).
@@
To further compare different centrality measures, we ran additional experiments on the Peptides and MolHIV datasets. We experimented in particular, with the Betweenness Centrality (BC), the Katz Index (KI) and the Subgraph Centrality (SC). We see from Table 7 that the performances achieved by different centrality measures are not dramatically different from each other, with those by the KI and SC being closer. In fact, centrality measures often exhibit a degree of correlation with each other, especially if from the same family, as it is the case of the walk-based KI and SC (see Estrada & Rodriguez-Velazquez (2005) and Table 6). It is also worth noting that Subgraph Centrality can be more efficient to calculate than these other centrality measures using the Networkx library (see Table 8).
@@
E.2 Comparison between Centrality-based Structural Encodings and RWSE
Here we aim to outline some of the similarities and differences between our Centrality structural Encoding (CSE) defined in Eq. 4 and the Random-Walk Structural Encoding (RWSE) introduced in (Dwivedi et al., 2021). The RWSE uses the diagonal of the $k$-step random-walk matrix defined in Eq. 44 defined as:
@@ E.2 Comparison between Centrality-based Structural Encodings and RWSE
These terms show similarity to our CSE as it also stores powers of the diagonal of the adjacency matrix, but it has a different normalization term that depends on the degree. In Proposition 1, we show that using an MPNN with CSE can compute the the probability, for all possible walks departing from a node, that a walk will lead back to the start. RWSE structural encodings are subtly different in that they compute the landing probability of a random walk from a node to itself. In this case, rather than weighting all possible walks equally, the walks are weighted by the degrees of the nodes along the walk; a walk where there are fewer alternative routes (other nodes) for the RW is more likely to occur.
@@
To answer (ii) and highlight the benefit of sampling based on CSE over using RWSE, we compared using the sum of these different encodings to sample the subgraphs in the counting substructures experiment. From Fig. 7, we see that our sampling method is better for counting all substructures and for all the number of samples in comparison to RWSE sampling.
In conclusion, Section F.2 shows that our CSE is better for sampling subgraphs and Table 9 shows that CSE is competitive when purely used as a Structural Encoding. Therefore, it is well motivated to use the CSE for our hybrid method where we need an SE and to use it as a sampling method. Future work could consider further understanding the expressivity differences between these SEs and the role of the normalization factor.
@@
E.3 Further examining the effect of CSEs
@@
Appendix F Experimental Details
@@ F.1 Architectural form
F.2 Synthetic Experimental Details
F.2.1 Dataset Generation
For the synthetic counting substructures experiment, we generated a dataset of random unattributed graphs in a similar manner to (Chen et al., 2020). In their experiments, they generate 5000 random regular graphs denoted as $RG(m,d)$, where $m$ is the number of nodes in each graph and $d$ is the node degree. Random regular graphs with $m$ nodes and degree $d$ are sampled and then $m$ edges are randomly deleted. In their work, Chen et al. (Chen et al., 2020), uniformly sampled $(m,d)$ from ${(10,6),(15,6),(20,5),(30,5)}$. However, we want to test the effectiveness of our sampling approach for larger sizes of graphs. Additionally, we found that the number of substructures present in the graph was related to the graphs size (see Fig. 8). Therefore, we wanted to create a more challenging benchmark with larger graph sizes. Therefore, we set $(m,d)$ to be $(60,5)$ for all graphs.
@@ F.2.1 Dataset Generation
F.2.2 Synthetic model parameters
@@
F.3 Real-World Experimental Details
@@ F.3.1 Hardware
F.3.2 Dataset specific details
@@
Appendix G Additional Time Comparisons

### v2

@@ F.4 Comparison to Other Centrality Measures
To further assess the benefits of our specific centrality encoding for sampling subgraphs, we compared against using other centrality measures to sample subgraphs in the counting substructure task. We used Closeness centrality, Betweeness centrality, Pagerank centrality and Degree centrality as baselines using the Networkx library (Hagberg et al., 2008). From Figure 8, we can see that using any of the different centrality methods performs better than random sampling across all substructures and number of samples. We also find that the Subgraph Centrality which we use, outperforms all other approaches in counting $3$ and $4$-cycles for any number of samples and in counting $3$ and $4$-paths when number of samples $\geq 5$. For $4$-cycles and other substructures, we find that Subgraph centrality is best, followed by Degree and Pagerank centrality and then Closeness and Betweenness centralities perform the worst of these centrality measures. This ranking of performance is aligned with how correlated these substructures are with the Subgaph Centrality on these synthetic graphs (as shown in Table 13).
@@
To further compare different centrality measures, we ran additional experiments on the Peptides and MolHIV datasets. We experimented in particular, with the Betweenness Centrality (BC), the Katz Index (KI) and the Subgraph Centrality (SC). We see from Table 14 that the performances achieved by different centrality measures are not dramatically different from each other, with those by the KI and SC being closer. In fact, centrality measures often exhibit a degree of correlation with each other, especially if from the same family, as it is the case of the walk-based KI and SC (see (Estrada & Rodriguez-Velazquez, 2005) and Table 13). It is also worth noting that Subgraph Centrality can be more efficient to calculate than these other centrality measures using the Networkx library (see Table 15).
@@
F.5 Comparison between Centrality-Based Structural Encodings and RWSE
Here we aim to outline some of the similarities and differences between our Centrality structural Encoding (CSE) defined in Equation 4 and the Random-Walk Structural Encoding (RWSE) introduced in (Dwivedi et al., 2021). The RWSE uses the diagonal of the $k$-step random-walk matrix defined in Equation 44 defined as:
@@ F.5 Comparison between Centrality-Based Structural Encodings and RWSE
These terms show similarity to our CSE as it also stores powers of the diagonal of the adjacency matrix, but it has a different normalization term that depends on the degree. In Proposition D.1, we show that using an MPNN with CSE can compute the the probability, for all possible walks departing from a node, that a walk will lead back to the start. RWSE structural encodings are subtly different in that they compute the landing probability of a random walk from a node to itself. In this case, rather than weighting all possible walks equally, the walks are weighted by the degrees of the nodes along the walk; a walk where there are fewer alternative routes (other nodes) for the RW is more likely to occur.
@@
To answer (ii) and highlight the benefit of sampling based on CSE over using RWSE, we compared using the sum of these different encodings to sample the subgraphs in the counting substructures experiment. From Figure 9, we see that our sampling method is better for counting all substructures and for all the number of samples in comparison to RWSE sampling.
In conclusion, Section G.2 shows that our CSE is better for sampling subgraphs and Table 16 shows that CSE is competitive when purely used as a Structural Encoding. Therefore, it is well motivated to use the CSE for our hybrid method where we need an SE and to use it as a sampling method. Future work could consider further understanding the expressivity differences between these SEs and the role of the normalization factor.
@@
F.6 Further Examining the Effect of CSEs
@@
Appendix G Experimental Details
@@ G.1 Architectural Form
G.2 Synthetic Experimental Details
G.2.1 Dataset Generation
For the synthetic counting substructures experiment, we generated a dataset of random unattributed graphs in a similar manner to (Chen et al., 2020). In their experiments, they generate 5000 random regular graphs denoted as $RG(m,d)$, where $m$ is the number of nodes in each graph and $d$ is the node degree. Random regular graphs with $m$ nodes and degree $d$ are sampled and then $m$ edges are randomly deleted. In their work, Chen et al. (Chen et al., 2020), uniformly sampled $(m,d)$ from ${(10,6),(15,6),(20,5),(30,5)}$. However, we want to test the effectiveness of our sampling approach for larger sizes of graphs. Additionally, we found that the number of substructures present in the graph was related to the graphs size (see Figure 10). Therefore, we wanted to create a more challenging benchmark with larger graph sizes. Therefore, we set $(m,d)$ to be $(60,5)$ for all graphs.
@@ G.2.1 Dataset Generation
G.2.2 Synthetic Model Parameters
@@
G.3 Real-World Experimental Details
@@ G.3.1 Hardware
G.3.2 Dataset Specific Details
@@
MalNet-Tiny (CC-BY license) This dataset comprises of function call graphs derived from Android APKs. The dataset used is a subset of MalNet and contains $5,000$ graphs of up to $5,000$ nodes, coming from benign software or $4$ types of malware. The task is to classify the type of software based on its structure.
Table 21: Best performing hyperparameters in Table 4.
Hyperparameter MalNet-Tiny
#Layers 5
#Layers readout 3
Hidden dim 96
Dropout 0.0
Subgraph pooling sum
Positional Encoding Steps 20
PE dim 28
#Parameters
REDDIT-BINARY (CC-BY license) The dataset consists of social networks posts are connected by an edge if the same user comments on both. This dataset is large in size, containing 232,965 posts with an average degree of 429. The task is to predict whether the graph belongs to a question/answer-based community or a discussion-based community. We used the evaluation procedure proposed in Xu et al. (2018), consisting of a 10-fold cross-validation and a metric with the best averaged validation accuracy across the folds.
Table 22: Best performing hyperparameters in Table 5.
Hyperparameter RDT-B
#Layers 2
#Layers readout 1
Hidden dim 300
Dropout 0
Subgraph pooling mean
Positional Encoding Steps 16
PE dim 16
#Parameters 274204
Appendix H Additional Time Comparisons